# Simulating Dual-Pixel Images From Ray Tracing For Depth Estimation

### v1

@@ 4 DfDP Model Based on Sdirt
As shown in Fig. 4(b), the input of the DfDP model is provided by the pixel-wise DP image rendering module of the Sdirt in real-time. In each training iteration, Sdirt feeds the DfDP model with simulated DP images of shape (B, C, H, W). Here, B is the batch size, C is the number channels, and H and W are the DP image height and width. The model outputs a predicted depth map of shape (B, 1, H, W). We use the L1 function as the loss:
$L=L_{1}(\hat{I_{D}},I_{D})$ (5)
We select [6] as the DfDP model and make reasonable adjustments to its cost volume step to accommodate the DP data. In binocular images, disparity formed by points at different depths is always in the same direction, thus the cost volume generation process in [6] is also unidirectional, as shown by the green connecting arrows Fig. 4(c). However, in DP images, disparity formed by object points before and after the focus distance is opposite in direction. Therefore, we add blue connecting arrows to extend the cost volume generation process, in order to enhance the DfDP model’s ability to capture reverse disparity.
@@ 5 Sdirt Representation Result
Figure 6: Qualitative results of simulated DP images. Evaluate the similarity between simulated (ours, COC, L2R [2], Modeling [26], and DDDNet [23]) and real F/4 defocused left DP images at two depths (0.5m and 2m). Compared to real F/4 defocused images, images simulated by other methods exhibit varying sizes of patterns, incomplete shapes, and texture shifts in different directions before and after the focus distance (1m). Our scheme produces the most realistic simulated images.
@@ 5.1 Implementation details
The Sdirt scheme employs a Canon RF50mm lens and a R6 Mark II camerabody as the our real camera. The simulated DP sensor has dimensions of $24mm\times 36mm$ and a resolution of $512\times 768$. We set the focus distance of both the simulated and real cameras to 1 meter, the valid imaging range to 0.5 to 20 meters, the F-number to F/4, and the ray set $A$ to containing 4096 rays.   When training the MLP, the kernel size $ks$ to 21 to fully display the DP PSF of any point within the valid imaging region. We train the network for 100,000 iterations, selecting 128 random points within the valid imaging region during each iteration. We conducted our experiments on a 12700K CPU and a 3090 GPU.
Following [39], we assume chromatic aberration is well corrected compared to other aberrations and defocus effects, allowing us to use a 550 nm wavelength to reduce ray tracing costs. For details on obtaining the structure of the DP composite pixel, please refer to the Supplementary.
@@ 5.2 Comparison methods
For comparison, we select all existing model-based DP PSF simulators: Modeling, L2R [2], DDDNet [23], and added a COC simulator. Although [24] is a more recent work, its simulator fully reuses DDDNet [23], so we do not include it for comparison. Calibration-based simulators and constraint-training methods are data-driven and require real DP-depth paired data, making them fundamentally different from model-based simulators, so no comparison is made. We directly used the source code of these DP PSF simulators [26, 2], and replicated the simulator [23] that only provided executable files. To ensure fairness, we set the focus distance, F-number, and sensor size in their simulator to be consistent with ours.
5.3 Evaluation on DP PSF simulator
To facilitate the observation of aberration and phase information, we temporarily set the F-numbers of both the simulated and real cameras to F/1.8. We capture a screen displaying a bright spot to obtain the real DP PSF.
As shown in Fig. 5, we evaluate the DP PSF ($PSF_{L}$ on left, $PSF_{R}$ on right) of spatial object points $p$ at various depths and positions within the valid imaging region. Observing the real DP PSFs, it can be found that defocus (focus distance is 1m) causes a significant phase difference between the real $PSF_{L}$ and $PSF_{R}$, and the phase difference at depths of 0.5m and 1.5m is opposite. When $p$ is on the optical axis, all other simulators fail to provide accurate DP PSF results close to reality. Farther from the optical axis, real $PSF_{L}$ and $PSF_{R}$ become asymmetric, with less significant phase differences and more off-axis aberrations. The differences between DP PSFs provided by other simulators and real DP-PSFs become more apparent. In contrast, our ray-traced simulator predicts accurate results at all depths and positions.
We present quantitative results in Tab. 1. Specifically, we choose the Normalized Squared Difference (NSD) method and the Normalized Cross-Correlation (NCC) method from OpenCV to evaluate the error and similarity between the real and simulated DP PSFs. We sampled 50 DP PSFs at 0.5m and 1.5m within the valid imaging region. Since the DP PSF is rotationally symmetric around the optical axis, we performed uniform sampling in the first quadrant. Our method achieves the highest similarity (0.915 NCC) and lowest error (0.133 NSD), demonstrating that we have addressed the domain gap between simulated and real DP PSFs caused by lens aberrations and sensor phase splitting.
Table 1: Quantitative results of DP PSF simulators. The NCC and NSD between the simulated and real DP PSFs at 50 points were evaluated. Our method achieved the best similarity and the lowest error.
@@ 5.3 Evaluation on DP PSF simulator
In subsequent experiments, the F-number will be adjusted back to F/4 due to the large blur kernel size caused by defocus at F/1.8, which resulted in GPU memory shortages.
5.4 Evaluation on simulated DP image
When comparing simulated and real DP images, to explore the degradation of simulated DP images caused by defocus effects at different depths, we captured richly textured planar images from 0.5m to infinity, perpendicular to the camera’s optical axis, with denser sampling at shallower depths of field. This resulted in 56 scenes with varying depths, including 11 with infinite depth. For each scene, real DP images were captured at F/4 and F/20. The F/20 image served as the AiF RGB image, replicating its capture-time depth information into the depth map. Then, we input the F/20 RGB-depth paired images into DP simulators to obtain the simulated F/4 defocused DP image.
As shown in Fig. 6, we compare the qualitative analysis results of the defocused left DP images at depths of 0.5m and 2m. Due to errors in simulating the DP PSF, other methods result in simulated images with textures shifted to the right at a depth of 0.5m and to the left at a depth of 2m, compared to the real left DP image. Moreover, other methods show different image texture sizes and incomplete shapes. However, the DP images simulated by our Sdirt method show minimal deviation in image textures.
Figure 7: Quantitative results of simulated DP images. Evaluate the similarity of 56 real and simulated (ours, COC, L2R [2], Modeling [26] and DDDNet [23]) planar scene F/4 DP images at different depths using PSNR ($\uparrow$) and SSIM ($\uparrow$) metrics. As the depth of simulated images deviates further from the focus distance (1m), the less realistic other methods become, whereas our scheme maintains the highest accuracy across all depths.
Figure 8: Qualitative results of depth estimation. Evaluate DfDP models by with COC, L2R [2], Modeling [26], DDDNet [23] and our Sdirt at four casual scenes. Each result image is decorated with a color bar in meters. Their depth estimation results show partial accuracy in relative positional relationships but large absolute positional errors. Our depth estimation results, however, demonstrate accurate in both relative and absolute positions with minimal errors. (Best viewed in colour and enlarge on screen.)
As shown in Fig. 7, which presents quantitative analysis results for all scenes, we assess the similarity between simulated and real images using PSNR and SSIM metrics. As the simulated image’s depth moves further away from the focus distance (1m), the neglect of aberrations and phase splitting in other methods leads to a more significant decrease in their metrics. In contrast, our Sdirt method achieves the best results at all depths, with average PSNR/SSIM values of 37.1982/0.9845. These values indicate that our simulated DP images are very realistic.
@@ 6.1 Implementation details
To train the DfDP model, we simulate F/4 DP images using the Sdirt based on RGBD datasets [30, 21] . NYU Depth dataset [30] is parsed by [12] and contains 50,688 indoor scenes. We use AdamW optimizer [19] and CosineAnnealing scheduler [20] with an initial learning rate of $1\times 10^{-4}$. The training batch size is 4, and we train for 50 epochs. For each epoch, 2000 DP-depth paired data are randomly selected for training. After training, we directly evaluate the model’s generalization on the real-world test set without any fine-tuning.
6.2 A new test set DP119
Compared to previous simulation methods, Sdirt requires more physical priors. As shown in Tab. 2, there is no public dataset that meets the following conditions: 1. Real DP-depth paired data for simulating DP images. 2. Known lens structure for ray tracing. 3. A fixed focus distance, as changing it alters the DP PSF of the same object point. Therefore, we collected the DP119 dataset, which consists of 45 planar scenes, 44 box scenes, and 30 casual scenes, totaling 119 scenes.
@@ 6.2 A new test set DP119
The casual scenes represent common scenarios, suitable for evaluating the robustness of simulation models. The planar and box scenes are richly textured and do not contain textureless areas, which helps the model utilize aberration and phase difference cues for depth estimation. Textureless areas, even when defocused, do not introduce any aberrations or phase differences, making it difficult to estimate depth and interfering with the evaluation. Therefore, planar and box scenes are ideal for assessing whether the model addresses the domain gap between simulated and real DP images. Real DP images are captured by the Canon RF50mm lens with a R6 Mark II at F/4 and 1m focus. GT depth maps for planar scenes are created by the depths at the capture time, while those for box and casual scenes are obtained from LiDAR scans. For more information about the DP119 test set, please refer to the Supplementary.
@@ 6.3 Evaluation
We present qualitative and quantitative results for the DfDP model trained by COC, L2R [2], Modeling [26], DDDNet [23] and our Sdirt. We deploy their DfDP model with their DP PSF simulator source code, and all methods use the same depth estimation network structure as ours for a fair comparison.
Figure 8 shows the depth estimation results for four casual scenes, with a color bar in meters added to decorate the images. It can be seen that although other methods provide relative positional information in the center region, they fail to correctly estimate relative positional relationships in the edge regions, and both the center and edge regions have large absolute positional errors. This is because the DP PSFs predicted by these methods exhibit phase symmetry and shift-invariance, while real DP PSFs exhibit significant phase asymmetry and aberrations (which do not have shift-invariance), causing their methods to deviate from the real DP PSF characteristics, resulting in large errors, especially in the edge regions where the errors are most severe. The results of L2R [2] and Modeling [26] are similar because the peak positions of their DP PSFs are quite close. In contrast, our method provides accurate relative positional information across the entire image area with minimal absolute positional errors, as our DP PSF is highly consistent with the real DP PSF in both the center and edge regions.
Table 3: Quantitative results of depth estimation. The DfDP model trained based on the Sdirt scheme achieves the best performance across most scenarios and evaluation metrics.
@@
We evaluate all models on the DP119 test set using the following metrics to assess the quantitative results: Mean Absolute Error (MAE), Mean Squared Error (MSE), Absolute Relative Error (Abs.r.), Squared Relative Error (Sq.r.), accuracy with $\delta<1.25$ (Acc-1) and accuracy with $\delta<1.25^{2}$ (Acc-2). Since the planar and box scenes are richly textured and do not have interference from textureless areas, the simulation models can rely on abundant aberration and phase difference cues for depth estimation. As shown in Tab. 3, in the simplest planar scene, our model achieves the best results across all metrics (0.9849 Acc-1). In the box scenes, all models experience a performance drop compared to the planar scenes, but we still achieve the optimal metrics (0.9474 Acc-1). These results demonstrate that our model is highly realistic and has effectively addressed the domain gap between simulated and real DP images. Furthermore, we test the robustness of all models on the casual scenes. Due to the textureless areas in the casual scenes, the performance of all models significantly degrades. However, our model still achieves the best metrics (0.8236 Acc-1), far surpassing the second-best model (0.3318 Acc-1). This further demonstrates that our model is the most realistic and has the best generalization performance. For more sample results from the DP119 test set, please refer to the Supplementary.
@@ 7 Conclusion and Discussion
Simulated DP images can address the scarcity of DP-depth paired data but face a domain gap between simulated and real DP data. In this work, we propose Sdirt to bridge this domain gap. Specifically, we calculate the DP PSF for points in object space using ray tracing, and leverage a network to predict it. Then, we render DP images based on the predicted DP-PSFs. Experimental results show that the proposed Sdirt scheme can simulate more realistic DP data. Moreover, depth estimation models trained based on Sdirt generalize better on real DP images.
We believe that the Sdirt scheme is not limited to DfDP tasks, it can provide extra depth information for any task with known optical imaging system parameters, promising significant applications in scenarios such as smartphones, automobiles, and microscopes in the future.   However, Sdirt is only applicable to cameras equipped with a fixed-focus lens (the structure must be open) and a DP sensor (DP images must be available). Currently, only Canon (5D4, R series) meets these requirements. To further expand the application of Sdirt, more camera and smartphone manufacturers need to make these data accessible.
@@
8 Sdirt Details
Table 4: Canon RF50mm F/1.8 lens data. The reference lens used in the paper.
@@ 8 Sdirt Details
Figure 9: The 2D lens structure with ray paths. (a) Canon RF50mm F/1.8 lens. (b) Canon RF35mm F/1.8 lens.
Table 5: Canon RF35mm F/1.8 lens data. The lens used in detecting DP composite pixel structure parameters.
@@
8.1 Structure parameters of DP composite pixels
We select the Canon RF50mm F/1.8 lens and Canon R6 Mark II camera body as our reference lens and DP sensor. Since the lens parameters are known, with detailed lens data presented in Tab. 4 and the 2D lens structure with ray paths illustrated in Fig. 9(a), it is feasible to perform ray tracing on the lens. However, performing ray tracing on the DP sensor is limited by camera manufacturer’s nondisclosure of microlens and sub-pixel structure parameters. Therefore, as described in Sec. 3.1 (Ray-traced DP PSF simulator) and Fig. 3(a) (DP composite pixel structure) of the main paper, we simplify the DP composite pixel structure by modeling the microlens as a thin lens with a radius of $r$ and a focal length of $f$, set $h$ to the distance between the sub-pixel and the microlens, set $w$ to the sub-pixel width, and set $ps$ to the size of the DP composite pixel.
We assign a set of possible values to each structure parameter of the DP composite pixel and identify the optimal parameter combination through grid searching across these parameters. Instead of conducting search experiments directly on the reference lens, we replace it with the Canon RF35mm F/1.8 lens to probe the structure parameters, ensuring isolation between the reference lens and the sensor. Detailed lens data for the Canon RF35mm F/1.8 lens are provided in Tab. 5, and its 2D lens structure with ray paths is illustrated in Fig. 9(b). In the valid imaging region (Fig. 4(a) in the main paper), we select 5 object points at different positions and capture real DP PSFs through the camera. By evaluating the error and similarity between the real DP PSF and the set of DP PSFs simulated with all structure parameters, we determine an optimal parameter combination:
$\begin{cases}h&=0.78*ps\\ f&=1.44*ps\\ w&=0.60*ps\\ r&=0.50*ps\end{cases}$ (6)
@@ 8.1 Structure parameters of DP composite pixels
Figure 10: Error and similarity matching results of structure parameters. In the matching results, the horizontal axis represents structural parameters (multiples of pixel size). The left vertical axis represents the NSD matching result between real and simulated DP PSFs, while the right vertical axis represents the NCC matching result. To facilitate display, when showing the matching result for any one parameter, we fix the other parameters at their optimal values according to to Eq. 6. Moreover, we use white dashed lines to mark the optimal value for each parameter.
@@
Specifically, we choose the Normalized Squared Difference (NSD) method and the Normalized Cross-Correlation (NCC) method from OpenCV to evaluate the error and similarity between the real and simulated DP PSFs. During the search experiments, we set the F-number to F/4.0, set the DP PSF kernel size to 35, and set the focal distance to infinity. As shown in Fig. 10, we present the results of the error and similarity matching. For ease of presentation, when displaying the matching results for any one parameter, we fix the other parameters at their optimal values according to Eq. 6.
@@
Figure 11: Qualitative analysis results of DP PSF for real and ray tracing. We select 5 points within the valid imaging region, and only their x-coordinates increased sequentially, corresponding to (a) - (e). We provide not only the comparison results of real and ray-traced DP PSF ($PSF_{L}$ on the left, $PSF_{R}$ on the right) for these 5 points. We also present the pixel distribution curves for the central row. As the object point $p$ moves farther away from the optical axis, the real $PSF_{L}$ and $PSF_{R}$ become more asymmetric. Our simulated DP PSF, obtained through ray tracing, aligns well with the real DP PSF.

### v2

@@ 4 DfDP Model Based on Sdirt
As shown in Fig.˜4(c), the input to the DfDP model is provided in real time by the pixel-wise DP image rendering module of Sdirt. At each training iteration, Sdirt feeds the DfDP model with simulated DP images of shape $(B,6,H,W)$, where $B$ is the batch size, the number of channels is 6, and $H$, $W$ are the DP image height and width. The model outputs the predicted depth maps ($\hat{I}_{D}$) of shape $(B,1,H,W)$. We apply the L1 loss to supervise $\hat{I}_{D}$:
$Loss=L_{1}(\hat{I_{D}},I_{D})$ (5)
We select [6] as the DfDP model and make reasonable adjustments to its cost volume step to accommodate the DP data. In binocular images, disparity from points at different depths always has the same direction, thus the cost volume generation process in [6] is also unidirectional, as shown by the green arrows in Fig.˜4(d). However, in DP images, disparity from object points before and after the focus distance is opposite in direction. Therefore, we add blue arrows to extend the cost volume generation, in order to enhance the DfDP model’s ability to capture reverse disparity.
@@ 5 Sdirt Representation Result
Figure 6: Qualitative results of simulated DP images. Evaluate the similarity between simulated (ours, CoC, L2R [3], Modeling [25], and DDDNet [23]) and real F/4 defocused left DP images at two depths (0.5 m and 2 m). Compared to real F/4 defocused images, images simulated by other methods exhibit varying sizes of patterns, incomplete shapes, and texture shifts in different directions before and after the focus distance (1 m). Our method produces the most realistic simulated images.
@@ 5.1 Implementation details
Sdirt employs a Canon RF50mm lens and an R6 Mark II camera body as the real camera. The simulated DP sensor has dimensions of $24\,\text{mm}\times 36\,\text{mm}$ with a resolution of $512\times 768$. We set the focus distance of both the simulated and real cameras to 1 m, the valid imaging range to 0.5–20 m, the F-number to F/4, and the ray set $A$ to contain 4096 rays. During MLP training, $ks$ is set to 21 to fully display the DP PSF of any point within the valid imaging region. We train the MLP for 100,000 iterations, selecting 128 random points within the valid imaging region during each iteration. We conducted our experiments on a 12700K CPU and a 3090 GPU.
Following [39], we assume chromatic aberration is well corrected compared to other aberrations and defocus effects, allowing us to use a 550 nm wavelength to reduce ray tracing costs. For details on obtaining the structural parameters of the DP pixel, please refer to the Supplementary.
@@ 5.2 Comparison methods
For comparison, we select all existing model-based DP PSF simulators: L2R [3], Modeling [25], DDDNet [23], and an added CoC simulator. Although [24] is a more recent work, its simulator fully reuses DDDNet [23], so we do not include it for comparison. Calibration-based simulators and constraint-training methods are data-driven and require real DP-depth paired data, making them fundamentally different from model-based simulators, so no comparison is made. We directly used the source code of these DP PSF simulators [25, 3], and re-implemented the simulator [23] which only provided executable files. To ensure fairness, we set the focus distance, F-number, and sensor size in each simulator to be consistent with ours.
5.3 Evaluation on simulated DP PSFs
To facilitate observation of aberration and phase information, we temporarily set the F-numbers of both the simulated and real cameras to F/1.8. We capture a OLED screen displaying a bright pixel to obtain the real DP PSFs.
As shown in Fig.˜5, we evaluate the DP PSF ($PSF_{L}$ on the left, $PSF_{R}$ on the right) of object points $p$ at various depths and positions within the valid imaging region. Observing the real DP PSFs, it can be found that defocus (with focus distance of 1 m) causes a significant phase difference between the real $PSF_{L}$ and $PSF_{R}$, and the phase difference at depths of 0.5 m and 1.5 m is opposite. When $p$ is on the optical axis, all other simulators fail to provide DP PSF results close to the real ones. Farther from the optical axis, the real $PSF_{L}$ and $PSF_{R}$ become phase asymmetric, with smaller phase differences and more off-axis aberrations. The differences between DP PSFs from other simulators and the real ones become more apparent. In contrast, our ray-traced simulator predicts accurate results at all depths and positions.
We present the quantitative results in Tab.˜1, using OpenCV’s normalized squared difference (NSD) and normalized cross-correlation (NCC) to quantify the error and similarity between the real and simulated DP PSFs. We sampled 50 DP PSFs at 0.5 m and 1.5 m within the valid imaging region. Since the real DP PSF has flip symmetry (not phase symmetry) about both the $x$ and $y$ axes, we performed uniform sampling in the first quadrant. Our method achieves the highest similarity (0.915 NCC) and lowest error (0.133 NSD), demonstrating its effectiveness in bridging the domain gap between simulated and real DP PSFs introduced by lens aberrations and sensor phase splitting.
Table 1: Quantitative results of simulated DP PSFs. The NCC and NSD metrics between simulated and real DP PSFs at 50 points were evaluated. Our method achieved the highest similarity and lowest error.
@@ 5.3 Evaluation on simulated DP PSFs
In all subsequent experiments, the F-number is adjusted back to F/4 due to the large blur kernel size caused by defocus at F/1.8, which resulted in GPU memory shortages.
5.4 Evaluation on simulated DP images
When comparing simulated and real DP images, to assess the degradation of simulated DP images caused by defocus effects at different depths, we captured richly textured planar scenes from 0.5 m to infinity, perpendicular to the optical axis. Shallower depths were sampled more densely to better capture defocus variations. This resulted in 56 scenes with varying depths, including 11 infinite-depth scenes. For each scene, real DP images were captured at F/4 and F/20. The F/20 image served as the AiF RGB image, and its capture-time depth was assigned to the entire depth map. The F/20 RGB-depth pairs were then input into DP simulators to generate the simulated F/4 defocused DP images.
As shown in Fig.˜S5, we compare the qualitative results of the defocused left DP images at depths of 0.5 m and 2 m. Due to errors in simulating the DP PSF, other methods produce simulated images with textures shifted to the right at 0.5 m depth and to the left at 2 m depth, compared to the real left DP image. Moreover, other methods show variations in image texture sizes and incomplete shapes. However, the DP images simulated by our Sdirt method show minimal deviation in texture details.
Figure 7: Quantitative results of simulated DP images. Evaluate the similarity of 56 real and simulated (ours, CoC, L2R [3], Modeling [25], and DDDNet [23]) planar scene F/4 DP images at different depths using PSNR ($\uparrow$) and SSIM ($\uparrow$) metrics. As the depth of simulated images deviates further from the focus distance (1 m), the less realistic the other methods become, whereas our scheme maintains the highest accuracy across all depths.
Figure 8: Qualitative results of absolute depth estimation on DP119. Evaluate DfDP models with CoC, L2R [3], Modeling [25], DDDNet [23] and our Sdirt at four casual scenes. Each result image includes a color bar showing depth in meters. Their depth estimation results show partial accuracy in relative positional relationships but large absolute positional errors. Our depth estimation results, however, demonstrate high accuracy in both relative and absolute positions with minimal errors. (Best viewed in color and enlarge on screen.)
As shown in Fig.˜7, which presents the quantitative results, we assess the similarity between simulated and real images using PSNR and SSIM. As the simulated image’s depth moves further away from the focus distance (1 m), the neglect of aberrations and phase splitting in other methods leads to a more significant decrease in their metrics. In contrast, our Sdirt method achieves the best results at all depths, with average PSNR/SSIM values of 37.1982/0.9845, indicating that our simulated DP images are highly realistic.
@@ 6.1 Implementation details
To train the DfDP model, we simulate F/4 DP images using Sdirt on RGBD datasets [30, 21]. The NYU Depth dataset [30], preprocessed by [12], contains 50,688 indoor scenes. We adopt the AdamW optimizer [19] and CosineAnnealing scheduler [20] with an initial learning rate of $1\times 10^{-4}$. The model is trained with a batch size of 4 for 50 epochs. In each epoch, 2,000 randomly selected DP-depth pairs are used for training. After training, the model is directly evaluated on a real-world test set without any fine-tuning to assess its generalization ability.
6.2 A new real-world test set DP119
Unlike previous simulation methods, Sdirt addresses data scarcity by simulating hardware, requiring consistent camera and setup during training and inference. However, as shown in Tab.˜2, no public dataset satisfies all of the following conditions: 1. Real DP-depth paired data for simulating DP images; 2. A known lens structure for ray tracing; 3. A fixed focus distance, since changing it alters the DP PSF of the same object point.   This makes evaluation on public datasets with unknown hardware impractical. We therefore collected the DP119 test set, which contains diverse test scenes.
@@ 6.2 A new real-world test set DP119
DP119 consists of 45 planar scenes, 44 box scenes, and 30 casual scenes, totaling 119 scenes. The casual scenes represent common scenarios, suitable for evaluating the robustness of simulation models. The planar and box scenes are richly textured, helping the model utilize aberration and phase cues for depth estimation. Textureless areas, even when defocused, lack aberrations or phase differences, making depth estimation difficult and evaluation unreliable. Thus, richly textured planar and box scenes are ideal for assessing domain gaps between simulated and real DP images. Real DP images are captured by the Canon RF50mm lens with a R6 Mark II at F/4 and 1 m focus. GT depth maps for planar scenes are created by the depths at the capture time, while those for box and casual scenes are obtained from LiDAR scans. For more information about the DP119 test set, please refer to the Supplementary.
@@ 6.3 Evaluation
We present qualitative and quantitative results for the DfDP model trained by CoC, L2R [3], Modeling [25], DDDNet [23], and our Sdirt. For a fair comparison, we deploy their DfDP models using their DP PSF simulator source code [3, 25] and re-implemented code [23], and have all models share the same depth estimation network architecture, initialize from a common CoC pre-trained checkpoint, and follow the same implementation details.
Figure˜8 shows depth estimation results for four casual scenes, with a color bar in meters for visualization. It can be seen that although other methods provide some relative positional information in the center region, they fail to correctly estimate relative positional relationships in the edge regions, and both the center and edge regions have large absolute positional errors. This is because the DP PSFs predicted by these methods exhibit phase symmetry and shift-invariance, whereas real DP PSFs exhibit phase asymmetry, aberrations, and shift-variance. Their methods’ mismatch with real DP PSFs leads to significant depth estimation errors, especially near the edges. The results of L2R [3] and Modeling [25] are similar because the peak positions of their DP PSFs are quite close. In contrast, our method provides accurate relative positional information across the entire image area with minimal absolute positional errors, as our simulated DP PSFs are highly consistent with the real ones in both the center and edge regions.
Table 3: Quantitative results of absolute depth estimation on DP119. The DfDP model trained with our Sdirt achieves superior performance across most scenarios and evaluation metrics.
@@
We evaluate all models on the DP119 test set using the following metrics to assess depth estimation performance: mean absolute error (MAE), mean squared error (MSE), absolute relative error (Abs.r.), squared relative error (Sq.r.), accuracy with $\delta<1.25$ (Acc-1) and accuracy with $\delta<1.25^{2}$ (Acc-2). Since the planar and box scenes are richly textured and do not have interference from textureless areas, the simulation models can rely on abundant aberration and phase difference cues for depth estimation. As shown in Tab.˜3, in the simplest planar scene, our model achieves the best results across all metrics (0.9849 Acc-1). In the box scenes, all models experience a performance drop compared to the planar scenes, but we still achieve the optimal metrics (0.9474 Acc-1). These results demonstrate that our model is highly realistic and effectively bridges the domain gap between simulated and real DP images. Furthermore, we test the robustness of all models on the casual scenes. Due to the textureless areas in the casual scenes, the performance of all models significantly degrades. However, our model still achieves the best metrics (0.8236 Acc-1), far surpassing the second-best model (0.3318 Acc-1). This further demonstrates that our model is the most realistic and has the best generalization performance. For more sample results from the DP119 test set, please refer to the Supplementary.
@@ 7 Conclusion and Discussion
Simulated DP images help address the scarcity of DP-depth paired data but face a domain gap between simulated and real DP data. In this work, we propose a novel simulation framework called Sdirt to bridge this gap. Specifically, we calculate the DP PSF for points in object space using ray tracing, and employ a network to predict them. Then, we render DP images based on the predicted DP PSFs. Experimental results show that the proposed Sdirt can simulate more realistic DP data. Moreover, depth estimation models trained based on Sdirt generalize better to real DP images.
We believe that Sdirt is not limited to DfDP tasks, it can provide extra depth information for any task with known optical imaging system parameters, promising significant applications in scenarios such as smartphones, automobiles, and microscopes in the future.   However, Sdirt is only applicable to cameras equipped with a fixed-focus lens (the structure must be open) and a DP sensor (DP images must be available). Currently, only Canon (5D4, R series) meets these requirements. To further expand the application of Sdirt, more camera and smartphone manufacturers need to make these data accessible.
Acknowledgments
This work was supported by National Natural Science Foundation of China (Grant No. 32471146) and the project N20240194. The authors thank Echossom, Miya, and Xinge for valuable discussions and assistance.
@@ Acknowledgments
S1 Sdirt Details
Figure S1: The 2D lens layout with ray paths. (a) Canon RF50mm F/1.8 lens. (b) Canon RF35mm F/1.8 lens.
Table S1: Canon RF50mm F/1.8 lens data. The real reference lens used in the main paper.
@@
Table S2: Canon RF35mm F/1.8 lens data. The real isolation lens used for detecting DP pixel structural parameters.
@@
Figure S2: Error and similarity matching results of DP pixel structural parameters using the isolation lens. For each subplot, the horizontal axis represents the corresponding structural parameter, where the values are given in multiples of the pixel size. The left vertical axis represents the NSD matching result between real and simulated DP PSFs, while the right vertical axis represents the NCC matching result. To facilitate display, when showing the matching result for any one parameter, we fix the rest of the parameters at their optimal values according to Eq.˜S1. Moreover, we use white dashed lines to mark the optimal value for each parameter.
S1.1 Structural parameters of the DP pixel
We select the Canon RF50mm F/1.8 lens and Canon R6 Mark II camera body as our real reference lens and DP sensor in the main paper. Since the reference lens data are available, with detailed specifications presented in Tab.˜S1 and the 2D lens layout with ray paths shown in Fig.˜S1(a), it is feasible to perform ray tracing on the lens. In practice, many commercial lens data, including Canon lenses, are available on open-access websites such as the Optical Bench, which facilitates optical modeling. However, performing ray tracing on the DP sensor is limited by camera manufacturers’ nondisclosure of the structural parameters of the microlens and sub-pixel components within the DP pixel. Therefore, as described in Sec. 3.1 (Ray-traced DP PSF simulator) and Fig. 3(a) (DP pixel structure layout) of the main paper, we simplify the DP pixel structure by modeling the microlens as a thin lens with radius $r$ and focal length $f$, defining $h$ as the distance between the sub-pixel and the microlens, $w$ as the sub-pixel width, and $ps$ as the DP pixel size.
We assign a set of possible values to each structural parameter of the DP pixel and identify the optimal parameter combination through grid searching across these parameters. Instead of conducting search experiments directly on the reference lens, we employ the Canon RF35mm F/1.8 as an isolation lens to replace the reference lens for probing DP pixel structural parameters, thereby ensuring isolation between the reference lens and the sensor. Detailed lens data for the Canon RF35mm F/1.8 are provided in Tab.˜S2, and its 2D lens layout with ray paths is shown in Fig.˜S1(b). In the valid imaging region (Fig. 4(b) in the main paper), we select 5 object points at different positions and capture real DP PSFs through the employed camera. By evaluating the error and similarity between the real DP PSFs and the set of DP PSFs simulated with all DP pixel structural parameters, we determine an optimal parameter combination:
$\begin{cases}h&=0.78*ps\\ f&=1.44*ps\\ w&=0.30*ps\\ r&=0.50*ps\end{cases}$ (S1)
@@
Specifically, we choose the normalized squared difference (NSD) and normalized cross-correlation (NCC) methods from OpenCV to evaluate the error and similarity between the real and simulated DP PSFs. During the search experiments, we set the F-number to F/4.0, set the DP PSF kernel size to 35, and set the focal distance to infinity. As shown in Fig.˜S2, we present the results of the error and similarity matching. For ease of presentation, when displaying the matching results for any one parameter, we fix the other parameters at their optimal values according to Eq.˜S1.
@@
Figure S3: Qualitative results of DP PSF for real and ray tracing. We select 5 points within the valid imaging region, and only their x-coordinates increased sequentially, corresponding to (a) - (e). We provide not only the comparison results of real and ray-traced DP PSF ($PSF_{L}$ on the left, $PSF_{R}$ on the right) for these 5 points. We also present the pixel distribution curves for the central row. As the object point $p$ moves farther away from the optical axis, the real $PSF_{L}$ and $PSF_{R}$ become more phase asymmetric. Our simulated DP PSFs, obtained through ray tracing, align well with the real DP PSFs.